OpenStudy (anonymous):

if x equals -1, find the matching value for y. 9x+2y=-5

OpenStudy (anonymous):

9(-1)+2y=-5 -9+2y=-5 2y=-5+9 2y=4 y=2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome